### v1

# Generalizing Visual Question Answering from Synthetic to Human-Written Questions via a Chain of QA with a Large Language Model

Abstract
Visual question answering (VQA) is a task where an image is given, and a series of questions are asked about the image. To build an efficient VQA algorithm, a large amount of QA data is required which is very expensive. Generating synthetic QA pairs based on templates is a practical way to obtain data. However, VQA models trained on those data do not perform well on complex, human-written questions. To address this issue, we propose a new method called chain of QA for human-written questions (CoQAH). CoQAH utilizes a sequence of QA interactions between a large language model and a VQA model trained on synthetic data to reason and derive logical answers for human-written questions. We tested the effectiveness of CoQAH on two types of human-written VQA datasets for 3D-rendered and chest X-ray images and found that it achieved state-of-the-art accuracy in both types of data. Notably, CoQAH outperformed general vision-language models, VQA models, and medical foundation models with no finetuning.
1 Introduction
Visual question answering (VQA) aims to build an automated algorithm to answer a series of questions regarding a given image. This task has a wide range of potential applications, such as interpreting medical images Lin et al. (2023) and supporting visually impaired people Bigham et al. (2010).
Recently, general vision language models (VLMs) Alayrac et al. (2022)OpenAI (2023)Liu et al. (2023), trained using a massive amount of image and text data, have shown promising results in solving VQA tasks. However, their performance is limited when tested in VQA tasks of specific domains (e.g., VLM model trained on large amounts of medical images and captions on the web vs. VQA model specialized for chest X-rays).
Figure 1: (a) Human-written questions, compared to the fixed format of the template-based questions, include more complex free-form questions, such as ones requiring reasoning. (b), (c) Example cases where a template-based VQA model fails to answer human-written questions correctly.
This challenge highlights the need to train or finetune VLM models using VQA data in a particular domain. However, acquiring such data is challenging, requiring experts to design VQA datasets carefully without contradiction and logical errors. For instance, to create a VQA dataset of chest X-rays, radiologists must write QA pairs consistent with the given radiographs.
To address this challenge, a template-based approach, which automatically synthesizes QA pairs based on pre-defined templates, has been adopted. This approach produced high-quality VQA datasets with minimum errors (e.g., CLEVR Johnson et al. (2017a) and MIMIC-Diff-VQA Hu et al. (2023)). Based on these datasets, some studies Hu et al. (2023)Hudson and Manning (2018)Perez et al. (2018)Kamath et al. (2021) reported very high accuracy on internal testing (e.g., 99.7% in CLEVR Kamath et al. (2021)). However, they also found that those VQA models failed to answer human-written questions that deviated from the templates, such as questions requiring complex reasoning (e.g., what shape is the large yellow object? vs. what shape do the most object have in common? in Figure 1b) and, even paraphrased questions (e.g., where is the lesion? vs. which area is the lesion present in? in Figure 1c).
In this study, we propose a novel method called “chain of QA for human-written questions” (CoQAH) which can correctly answer complex, human-written questions beyond a fixed set of synthetic, template-based questions without the need for finetuning. In the core part of CoQAH, a large language model (LLM) sequentially asks template-based questions to a VQA model trained with synthetic data. This step-by-step QA process enables the LLM to collect valuable and accurate information about an image and reach a reasonable conclusion for a given question.
To check the effectiveness of CoQAH, we utilized two different types of template-based VQA datasets (CLEVR Johnson et al. (2017a) for 3D-rendered images and MIMIC-Diff-VQA Hu et al. (2023) for chest X-rays) for AI training. We then used multiple human-written VQA datasets (CLEVR-Human Johnson et al. (2017b), VQA-RAD Lau et al. (2018), and SLAKE Liu et al. (2021)) for AI testing. As a result, CoQAH achieved state-of-the-art performance on all the human-written VQA datasets, surpassing general VLMs and other VQA models.
Our contributions are summarized as follows: (1) we propose CoQAH, a new method that answers human-written questions through a step-by-step QA process between an LLM and a VQA model. (2) our method achieved state-of-the-art performance on multiple human-written VQA datasets, surpassing general VLMs and other VQA models. (3) we demonstrated the effectiveness of CoQAH on two types of VQA datasets, including 3D-rendered images (CLEVR and CLEVR-Human) and chest X-rays (MIMIC-Diff-VQA, VQA-RAD, and SLAKE).
2 Related Works
2.1 General Vision Language Models
A general VLM is an AI model that can understand images and texts, along with their relationships, in various tasks. One of the earliest approaches to creating a general VLM was to combine separate vision and text encoders, aligning features from both encoders Radford et al. (2021)Yu et al. (2022). Another approach utilized a pre-trained LLM, which can understand the meaning and context of texts. Tsimpoukelli et al. (2021) proposed training and combining a vision encoder with a frozen LLM. Similarly, Alayrac et al. (2022) proposed finetuning the connecting layers between an LLM and a vision encoder. More recently, Liu et al. (2023) proposed training VLMs using the instructions of various tasks to make a general-purpose assistant. These general VLMs can be applied to various vision-language tasks, but they often lack expertise in specific domains. CoQAH addresses this issue by employing a specialized VQA model (e.g., chest X-ray VQA) that interacts with an LLM.
2.2 Template-Based VQA
Creating VQA datasets manually is a time-consuming and expensive process. To overcome this challenge, researchers Johnson et al. (2017a)Hu et al. (2023) have attempted to synthesize QA pairs using pre-defined templates. For instance, one study Hu et al. (2023) introduced a VQA dataset for chest X-rays by extracting abnormal findings from radiologist’s reports. Similarly, another study Johnson et al. (2017a) created 3D-rendered images of various objects and generated complex template-based questions related to the visual features of those images. Several other studies Hudson and Manning (2018)Perez et al. (2018)Kamath et al. (2021)Hu et al. (2023) have used such datasets to evaluate and compare the visual understanding capabilities of AI models.
2.3 LLM-Aided VQA
Several studies Yang et al. (2022)Tiong et al. (2022)Guo et al. (2023)Zhou et al. (2023)You et al. (2023) have used an LLM to solve VQA problems. This is done by feeding the relevant information to the LLM, which then uses its high contextual understanding and reasoning ability to answer the question. For example, some researchers Yang et al. (2022)Tiong et al. (2022)Guo et al. (2023) have proposed feeding an image description from a captioning model to an LLM, while other Zhou et al. (2023) utilized VQA and captioning models together to give visual clues to an LLM. Recently, some researchers You et al. (2023) have suggested using interactions between an LLM and a VQA model to answer complex visual questions. CoQAH uses a template-based VQA model that is specialized to questions in specific domains, which makes it different from other methods above.
3 Method
Figure 2: An overview of the proposed CoQAH method. (a) An example of a task instruction is shown at the top. The figure on the left describes an overall interaction process between an LLM and a template-based VQA model to reach the final answer to the user question. On the right, the figure represents an example of dialogue between the two models. (b) The template format for the questions in the CLEVR dataset is described, which is composed of several different entities ($\langle question\rangle=\langle type\rangle+\langle object\rangle+\langle relation\rangle+\langle object\rangle$). For each entity, a few options are available to be selected (e.g., small or large or $\langle Empty\rangle$ for $\langle Size\rangle$ entity).
3.1 CoQAH
The CoQAH method aims to conclude a reasonable answer for a given complex, human-written question (i.e., user question; see Figure 2a) by employing synthetic, template-based VQA data only. To achieve this, it utilizes interactions between two sub-models, a template-based VQA model ($VQA_{t}$; see Supplementary Table 1 and 2 for the performance of VQA models) and a large language model ($LLM$).
Suppose a user question and an image from an human-written VQA dataset ($q_{h}\in Q_{h},i_{h}\in I_{h};Q_{h}$ for human-written questions and $I_{h}$ for images) are given; first, the $LLM$ is prompted with a task instruction ($\rho$; see Section 3.2 for detail) that directs the model to generate a template-based question that conforms to the question templates used to train the VQA model ($q_{t}^{1}=LLM(\rho);$ $q_{t}^{1}\in Q_{t}$ where $Q_{t}$ includes all possible template-based questions; see Section 3.2). Next, $VQA_{t}$ answers the question ($a_{t}^{1}=VQA_{t}(q_{t}^{1},i_{h})$). After that, all the previous dialogue, including the task instruction, template-based questions, and answers from $VQA_{t}$ $(\rho+q_{t}^{1}+a_{t}^{1})$, is again fed into $LLM$, generating a next template-based question ($q_{t}^{2}=LLM(\rho,q_{t}^{1},a_{t}^{1})$). This QA process is repeated until $LLM$ has sufficient information to reach the final answer for the user question or it reaches the maximum number of questions to be queried (twenty questions for CLEVR-Human and five questions for VQA-RAD and SLAKE; see Section 4.4). Note that $LLM$ cannot access the image, whereas $VQA_{t}$ perceive it.
3.2 Task Instruction
The task instruction $\rho$ is a detailed description of how $LLM$ should respond to and interact with $VQA_{t}$ (Figure 2a). The instruction starts with your goal is to answer this user question based on an image: followed by the user question. Then, we included a statement to declare that $LLM$ cannot see a given image (but you can’t access the image, and I can access the image.). After that, we asked $LLM$ to give a question that can only be generated based on a template format (you can ask me questions with these forms; see Section 3.3 for details). In the last part, we added some sentences to ask $LLM$ to conclude the answer for the user question whenever it is ready (if you can answer the user question, stop asking and give me an answer.).
3.3 Template-Based Questions from LLM
Figure 2b describes how we instructed $LLM$ to generate a template-based question for the CLEVR dataset Johnson et al. (2017a). We first let $LLM$ know the structure of the question, which is composed of several different entities ($\langle question\rangle=\langle type\rangle+\langle object\rangle+\langle relation\rangle+\langle object\rangle$ in Figure 2b). Subsequently, we asked $LLM$ to choose one of the options available for each entity (e.g., small or large or $\langle Empty\rangle$ for $\langle Size\rangle$ entity). The template format for MIMIC-Diff-VQA is shown in Supplementary Figure 1. The detailed description of the task instructions for both CLEVR and MIMIC-Diff-VQA is displayed in Supplementary Figure 2 and 3.
3.4 Existence and Uniqueness Handler
Even if $LLM$ successfully generated a template-based question, in some cases, we observed that there were still some questions that had logical errors (see Figure 3): questions that ask the property of an object that does not exist in the image (existence violation in Figure 3b) and ask the property of multiple objects with the same condition (uniqueness violation in Figure 3c). For these types of questions, $VQA_{t}$ provided incorrect answers unexpectedly.
To address this problem, we introduced an existence and uniqueness handler (EUH), which continually checks the violation of both conditions on a template-based question generated from $LLM$. As described in Algorithm 1, we first extracted the object entity from the question (i.e., $\langle object\rangle$ or $\langle object\rangle+\langle relation\rangle+\langle object\rangle$) and checked the existence of the object entities by querying to $VQA_{t}$ (is there a …? in Algorithm 1). When the existence of the object was confirmed, we continued to check the uniqueness by asking $VQA_{t}$ to count the number of objects (how many …? in Algorithm 1). Using EUH, we prevented $LLM$ from directly providing a wrong answer without considering logical contradiction.
Figure 3: Illustration of how the existence and uniqueness handler (EUH) can prevent an LLM from concluding an incorrect answer for a given user question. (a) An example case includes a question, an answer, and an image. (b) A dialogue between an LLM and a VQA model when the existence of an object is violated. With EUH, the VQA model checks the presence of an object and lets LLM know its existence. (c) A dialogue when the uniqueness of an object is violated. With EUH, the VQA model successfully lets LLM know the number of objects satisfying the same condition.
Algorithm 1 Existence and uniqueness handler
Input: Question from LLM $Q_{L}$, VQA model $VQA$
Output: Answer to $Q_{L}$ $A_{L}$
1:  $Extract\_Entity$: Extract (Object + Relation + Object) part from the question (Type + Object + Relation + Object)
2:  $E\leftarrow Extract\_Entity(Q_{L})$
3:  $Q_{E}\leftarrow$ “Is there a” $+$ $E$ $+$ “?”
4:  $A_{E}\leftarrow VQA(Q_{E})$
5:  if $A_{E}$ is “No” then
6:     $A_{L}\leftarrow$ “There is no” $+$ $E$
7:  else if $A_{E}$ is “Yes” then
8:     $Q_{U}\leftarrow$ “How many” $+$ $E$ $+$ “are there?”
9:     $A_{U}\leftarrow VQA(Q_{U})$
10:     if $A_{U}$ is not “1” then
11:        $A_{L}\leftarrow$ “There are” $+$ $A_{U}$ $+$ $E$
12:     else if $A_{U}$ is “1” then
13:        $A_{L}\leftarrow VQA(Q_{L})$
14:     end if
15:  end if
16:  Return $A_{L}$
4 Experiment Setup
4.1 Datasets
For datasets, we collected two types of images (and question-answer pairs), including 3D-rendered images and chest X-ray images, to show the effectiveness of CoQAH in different image types. We first trained VQA models using template-based datasets (CLEVR Johnson et al. (2017a) for 3D-rendered and MIMIC-Diff-VQA Hu et al. (2023) for chest X-ray) and externally validated the model’s performance on the human-written datasets (CLEVR-Human Johnson et al. (2017b) for 3D-rendered; VQA-RAD Lau et al. (2018) and SLAKE Liu et al. (2021) for chest X-ray).
CLEVR.
Johnson et al. (2017a) is a synthetic VQA dataset (70,000 image and QA pairs for training, 15,000 for validation, and 15,000 for testing). 3D objects of different sizes, colors, materials, and shapes were rendered for each image. Then, questions were generated based on a template-based algorithm, and the answers for those questions were automatically selected among a set of candidates (e.g., ”3”, ”blue”, ”cylinder”, etc.).
CLEVR-Human.
Johnson et al. (2017b) is a variant of the CLEVR dataset that replaced questions and answers in some images of CLEVR using human questioners (17,817 for training; 7,202 for validation). The questioners created new human-written questions without any format restriction, while the answers for the questions were chosen among the same candidates in CLEVR. Currently, a test set is not available. Therefore, we reported AI performance using the validation set.
MIMIC-Diff-VQA.
Hu et al. (2023) is a chest X-ray VQA dataset. QA pairs were generated using an automatic template algorithm based on radiologists’ reports (700,703 QA pairs from 164,324 reports and images; 80% for training, 10% for validation, and 10% for testing).
VQA-RAD and SLAKE.
Lau et al. (2018) and Liu et al. (2021) are medical image VQA datasets where clinical trainees (for VQA-RAD) or physicians (for SLAKE) wrote human-written questions and answers manually for each image. We used only the chest X-rays in those datasets (107 images for VQA-RAD and 179 images for SLAKE) for AI evaluation. In both datasets, answers are categorized as either closed-form or open-form: the answers of the closed-form questions were required to be one between two or three options (e.g., yes or no; 511 questions for VQA-RAD and 663 questions for SLAKE), whereas the answers of the open-form questions did not have any format (283 questions for VQA-RAD and 1,459 questions for SLAKE).
4.2 Metrics
[…]
4.3 Benchmarks
4.3.1 Benchmarks for CLEVR-Human
To benchmark the CLEVR-Human dataset, we employed three types of AI models, including general VLMs, VQA models trained with CLEVR (i.e., template-based VQA), and finetuned with CLEVR-Human (i.e., finetuned VQA). For the general VLMs, we utilized LLaVA Liu et al. (2023) (version 1.5; 7B and 13B parameters; source code and model weights available on GitHub; see task instruction in Supplementary Figure 4) and GPT-4-Vision OpenAI (2023) (version 1106-preview; API used; see task instruction in Supplementary Figure 5), testing them on the CLEVR-Human validation dataset. For the template-based VQA models, we benchmarked FiLM Perez et al. (2018), MAC Hudson and Manning (2018), MDETR Kamath et al. (2021), and our CoQAH. In CoQAH, we combined MDETR trained with CLEVR as a VQA model and GPT-4 as an LLM, measuring accuracy on the CLEVR-Human validation data. Finally, we compared the reported scores from the previous studies for the accuracy of the finetuned VQA models (FiLM, MAC, and MDETR) Perez et al. (2018)Hudson and Manning (2018)Kamath et al. (2021).
4.3.2 Benchmarks for VQA-RAD and SLAKE
[…]
4.3.3 Post-Processing of Answers
We observed that some of the answers the AI models provided frequently had the same meaning but different expressions (e.g., radiograph vs. x-ray, pa views vs. pa, right side vs. right, etc.), resulting in misleading scores. Therefore, we substituted those answers as standard forms using post-processing (see Supplementary Table 3). We applied this post-processing to all the AI models we tested for fair comparison.
4.4 Other Analyses
Since the CoQAH method utilizes an LLM to understand the context and finally answer the user question, we can ask the rationale of the final answer to the LLM (i.e., interpreting the AI model’s behavior), simply querying ’why?’. For a few complex questions, we asked the LLM to explain the reasons to check whether it had logically driven their conclusions.
In addition, to validate the effect of the maximum number of questions to be queried in CoQAH, we measured the accuracy of CoQAH on the CLEVR-Human validation dataset (5% used), changing that number from 0 to 30.
As an ablation study, we excluded EUH in CoQAH and measured the accuracy based on the CLEVR-Human validation data (5% used). To check the importance of the step-by-step QA process in CoQAH, we enforced the LLM to ask all the questions simultaneously (ask me up to 20 questions all at once), measuring the accuracy. Also, we let the LLM generate any form of questions during the chain of QA process by ablating the format description in the task instruction, also calculating accuracy.
Finally, we performed an error analysis by investigating the dialogues where CoQAH failed to answer correctly (5% CLEVR-Human validation used) and categorized them according to the reasons (e.g., incorrect answers from VQA model).
5 Results
Trained on Finetuned on Acc. on
CLEVR? CLEVR-Human? CLEVR-Human (%)
General VLMs
LLaVA (7B) Liu et al. (2023) No No   $*{}^{*}$
LLaVA (13B) Liu et al. (2023) No No   $*{}^{*}$
GPT-4-vision OpenAI (2023) No No   $*{}^{*}$
Template-Based VQA Models
FiLM Perez et al. (2018) Yes No 56.6
MAC Hudson and Manning (2018) Yes No 57.4
MDETR Kamath et al. (2021) Yes No 59.9 ($*{}^{*}$)
CoQAH Yes No   $*{}^{*}$
Finetuned VQA Models
FiLM Perez et al. (2018) Yes Yes 75.9
MAC Hudson and Manning (2018) Yes Yes 81.5
MDETR Kamath et al. (2021) Yes Yes 81.7
Table 1: Comparison of accuracy between the general VLMs, template-based VQA models, and finetuned VQA models. CoQAH achieved the highest accuracy among all the template-based models, although it did not utilize any data from CLEVR-Human for training. * indicates accuracy measured using the CLEVR-Human validation data.
VQA-RAD SLAKE
[…]
Closed-Form (%) on Open-Form Closed-Form (%) on Open-Form
Medical Foundation Models
BiomedGPT Zhang et al. (2023a) 43.2 0.150 - -
Med-Flamingo Moor et al. (2023) 46.4 0.184 33.2 0.140
MedVInT-TD Zhang et al. (2023b) 57.3 0.274 46.5 0.396
Template-Based VQA Models
OFA-MIMIC Wang et al. (2022) 59.5 0.095 69.4 0.097
CoQAH 67.5 0.302 73.9 0.425
[…]
5.1 Benchmarks on CLEVR-Human
Table 1 compares accuracy between the general VLMs, template-based, and finetuned VQA models. Overall, the general VLMs reported lower accuracy than the template-based VQA models, except for the latest GPT-4-vision (e.g., 60.1% for GPT-4-vision vs. 59.9% for MDETR). CoQAH achieved the highest accuracy with large gaps compared to all the general VLMs and template-based VQA models (e.g., 74.3% accuracy for CoQAH vs. 59.9% for MDETR). Although there are still some performance gaps between CoQAH and the finetuned models (e.g., 74.3% for CoQAH vs. 81.7% for finetuned MDTER), it is very difficult to obtain a large number of human-written QA pairs in general to finetune the model accordingly.
5.2 Benchmarks on VQA-RAD and SLAKE
[…]
5.3 Interpreting Reasons for Answers
Figure 4 shows some examples that we asked an LLM in CoQAH to interpret the reasons for answers. As shown in the examples, CoQAH successfully delineated the rationale behind the answers based on the information collected through the dialogues. For example, for the question of whether there are two cyan balls of the same size (Figure 4a), CoQAH reasonably explained the reason (one is large and one is small) for the answer (No) after checking the size of the cyan balls (Is there a large cyan ball? and Is there a small cyan ball?). More examples can be found in Supplementary Figure 8 (for 3D-rendered images) and 9 (for chest X-rays).
Figure 4: Example dialogues of CoQAH when we asked an LLM to interpret the reasons for the answers. In all cases, CoQAH successfully delineated the rationale behind the answers based on the information collected through the dialogues. For instance, for the question of Are the cyan balls are the same size?, CoQAH correctly explained the reason (one is large and one is small) for the response (No) based on the information from the two previous questions (Is there a large cyan ball? and Is there a small cyan ball?).
5.4 Effect of the Maximum Number of Questions
Table 3 shows the accuracy of CoQAH on the CLEVR-Human when we changed the maximum number of questions to be queried. The maximum accuracy was achieved when the maximum number of questions was 20, and even if we increased the number to 30, the accuracy was the same.
Maximum number of 0 5 10 20 30
questions to be queried
Acc.on 60 68.1 69.4 75.8 75.8
CLEVR-Human (%)
Table 3: Accuracy on CLEVR-Human validation dataset according to the changes of the maximum number of questions to be queried in CoQAH. We only used 5% of the validation data for the ablation study.
5.5 Ablation Study
The results of the ablation study are summarized in Table 4. When we ablated EUH in CoQAH, the accuracy on CLEVR-Human was slightly decreased (75.8% for CoQAH vs. 74.7% without EUH). When ablating the step-by-step QA process in CoQAH, the accuracy was degraded mainly (75.8% for CoQAH vs. 63.1% without the process), indicating the importance of this process to derive the answers correctly.Finally, when we ablated the template format description in the task instruction, the accuracy was dramatically degraded (75.8% for CoQAH vs. 35.8% without the format description). This means that the questions from the LLM must conform to the template format so the VQA model can understand those questions.
Acc. on
CLEVR-Human (%)
CoQAH 75.8
without EUH 74.7
without step-by-step QA process 63.1
without format description 35.6
Table 4: Comparison of accuracy on CLEVR-Human by ablating either EUH or step-by-step QA process. Ablating the QA process degraded the accuracy significantly. Ablating the template format description dramatically degraded the accuracy, meaning that the questions from the LLM must conform to the template format.
5.6 Error Analysis of CoQAH
The most significant portion of errors (48%) occurred when the LLM failed to ask proper questions to derive correct answers (e.g., Supplementary Figure. 8e). The second largest portion (25%) was due to incorrect responses from the VQA model during the QA process (e.g., Supplementary Figure. 8d). Some errors (11%) dedicated to the questions where the LLM was impossible to answer by querying template-based questions only (e.g., Supplementary Figure. 8f). Other errors (15%) included failures of following instructions, data format (e.g., The answer should be in {}), and so on.
6 Conclusion
In this study, we proposed a novel CoQAH methodology that answers human-written questions via a reasoning process based on a dialogue. Throughout the extensive experiments, we have shown that CoQAH achieved state-of-the-art performance on various datasets of different images (3D-rendered and X-ray) without finetuning using the human-written questions. We believe that this new method potentially promotes the wide applications of AI for visual question answering, improving the practicality of AI in real environment.
7 Limitation
First, we could only test a limited number of general VLM models (LLaVA Liu et al. (2023) and GPT-4-Vision OpenAI (2023)) for CLEVR-Human since, currently, many LLMs do not support taking an image as an input. Second, we reported the accuracy of the AI models we ran on CLEVR-Human validation data because the test set was unavailable. Third, even though we tried our best to optimize the task instructions for the general VLM and medical foundation models, those might not be the best due to the lack of methods to tune the instructions. Fourth, we only validated CoQAH using two types of images (3D-rendered and chest X-ray) because of a limited number of template-based datasets.
Ethical Statement
There are no ethical issues.
Acknowledgments
Acknowledgments
Supplementary Materials
Model Accuracy (%)
FiLM Perez et al. [2018] 97.7
MAC Hudson and Manning [2018] 98.9
MDETR Kamath et al. [2021] 99.7
Table 1: The performance of VQA models on the CLEVR testing dataset. Regardless of the models, the accuracy values were very high ($>$ 97%), meaning that those VQA models can precisely answer the template-based questions. We used the MDETR model as a template-based VQA model for CoQAH.
Model BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE CIDEr
EKIDG Hu et al. [2023] 0.624 0.541 0.477 0.422 0.337 0.645 1.893
OFA-MIMIC Wang et al. [2022] 0.662 0.588 0.528 0.473 0.362 0.716 2.332
[…]
Figure 1: Template format for the chest X-ray VQA dataset, MIMIC-Diff-VQA. A question is only composed of two parts, $\langle type\rangle$ and $\langle abnormality\rangle$. Similar to the case of the CLEVR dataset, we asked an LLM to select an option available for each $\langle type\rangle$ and $\langle abnormality\rangle$ to generate a template-based question.
Figure 2: Task instruction of CoQAH for CLEVR.
Figure 3: Task instruction of CoQAH for MIMIC-Diff-VQA.
Figure 4: Task instruction of LLaVA for CLEVR-Human. (a) In the first stage, we asked LLaVA to produce an answer as a single word or phrase. (b) In some cases, LLaVA provided answers of the same meaning but different expressions. Therefore, in the second stage, we corrected them by asking LLaVA to choose one of the candidates (e.g., round for the first answer vs. sphere for the corrected answer). Note that when we gave all options for the answer at once, the performance of LLaVA was primarily degraded.
Figure 5: Task instruction of GPT-4-Vision for CLEVR-Human. We provided a set of answers.
Figure 6: Task instruction of Med-Flamingo for VQA-RAD and SLAKE. We limited the maximum number of output words to one for the closed-form questions and five for the open-form questions.
Figure 7: Task instruction of MedVInT-TD for VQA-RAD and SLAKE. We limited the maximum number of output words to one for the closed-form questions and five for the open-form questions.
Synonym standardization for CLEVR-Human
Before standardization After standardization
tiny small
big large
matte rubber
metallic metal
shiny metal
block cube
ball sphere
Synonym standardization for VQA-RAD and SLAKE
Before standardization After standardization
x-ray xray
x ray xray
radiography xray
radiograph xray
cxr chest xray
pa view pa
ap view ap
right side right
left side left
Table 3: Table summarizing the answers of AI we substituted when calculating accuracy. We changed them because those words had the same meaning but different expressions, resulting in misleading performance metrics.
Figure 8: Example dialogues of CoQAH for CLEVR-Human. (a)-(c) The cases show that CoQAH correctly answered the questions and explained the reasons. (d) The answer was wrong because the template-based VQA model failed to give correct information. (e)-(f) CoQAH failed since the large language model could not ask relevant questions to derive correct answers.
Figure 9: Example dialogues of CoQAH for VQA-RAD and SLAKE. (a)-(c) CoQAH successfully answered the given user questions. (d) The VQA model failed to give the correct answer, and CoQAH failed accordingly. (e) The case shows that the large language model logically derived the answer even though the answer was incorrect. (f) Since the VQA model could not detect the lesion, CoQAH also gave the wrong answer.